T3=24 T6=60
FIND THE COMMON DIFF AND THE FIRST TERM (a)

Answer by Theo(13342)

without going through the formula, you can find the common difference in the following manner.

you are given that A3 = 24 and A6 = 60.
6 - 3 = 3
60 - 24 = 36
36/3 = 12

your common difference is 12.

your first term is A1.
A3 - A1 = 2
24 - 2*12 = 24 - 24 = 0
A1 = 0

A6 - A1 = 5
60 - 5*12 = 60 - 60 = 0
A1 = 0

the common difference is 12.
your first term is 0.

if you wish to use the formula to solve for this, do the followng.

the formula is An = A1 + (n-1) * d

this makes:
24 = A1 + 2 * d
60 = A1 + 5 * d

subtract the first equation from the first to get:
36 = 3 * d
solve for d to get d = 12.

go to either one of the formulas and replace d with 12.
24 = A1 + 24
solve for A1 to get A1 = 0

do the same thing with A6 and you get A1 = 0 as well, as you should, since A1 is the same in both formulas.